### v1

# Coning off totally geodesic boundary components of a hyperbolic manifold

Abstract.
Let $M$ be a compact hyperbolic manifold with totally geodesic boundary. If the injectivity radius of $\partial M$ is larger than an explicit function of the normal injectivity radius of $\partial M$, we show that there is a negatively curved metric on the space obtained by coning each boundary component of $M$ to a point. Moreover, we give explicit geometric conditions under which a locally convex subset of $M$ gives rise to a locally convex subset of the cone-off.
Group-theoretically, we conclude that the fundamental group of the cone-off is hyperbolic and the $\pi_{1}$-image of the coned-off locally convex subset is a quasi-convex subgroup.
1. Introduction
The idea of coning off ends or boundary components of an aspherical manifold to obtain an aspherical space goes back at least to a comment of Gromov in [Gro93, 7.A.VI]. This was fleshed out in the non-compact real hyperbolic case by Mosher and Sageev [MS97]. Mosher and Sageev begin with a hyperbolic manifold with cusps. Under the condition that disjoint cusp cross sections with large (greater than $\pi$) injectivity radius can be chosen, they show that a negatively curved space can be obtained by cutting off the cusps and replacing them with cones. The method is differential geometric, utilizing the same kind of warped product construction as in the Gromov-Thurston $2\pi$ theorem [BH96]. Modifying the construction, a wide variety of compact non-positively curved spaces can be obtained by replacing the cusps with mapping cylinders of fibrations [FM10].
In the current paper we consider a related but different starting point: a compact hyperbolic manifold with totally geodesic boundary. We give explicit geometric conditions under which a similar “coning” construction gives rise to a negatively curved space. Again, we need a lower bound on the injectivity radius of the boundary. However, this bound is not absolute (as in the $2\pi$ theorem) but depends on the size of a maximal collar neighborhood of the boundary.
We also give conditions of the same flavor under which a closed locally convex subset of the original manifold gives a locally convex (and hence $\pi_{1}$–injective) subset of the coned-off space. These results are used in work of the second author with Ruffoni to give examples of negatively curved three-dimensional pseudo-manifolds whose fundamental groups do not act properly on any CAT$(0)$ cube complex [MR].
As a consequence of our geometric construction, we obtain group-theoretic conclusions about the fundamental groups of the coned-off spaces and of the coned-off locally convex subsets. These are quantitative versions of results in “group-theoretic Dehn filling” [GM08, Osi07, GM21].
1.1. Statement of main theorems
First we say explicitly what we mean by coning off the boundary.
Definition 1.1.
Let $M$ be a compact manifold with boundary. The cone-off $\widehat{M}$ of $M$ is the space obtained by attaching to $M$ a cone on each boundary component. Formally we write
[…]
where we declare $x\sim(x\times 1)$ for any $x\in\partial M$ and $(x\times 0)\sim(y\times 0)$ whenever $x$ and $y$ lie in the same boundary component.
If $S\subset M$, we define $\widehat{S}\subset\widehat{M}$ to be the union of $S$ with the set of equivalence classes of points of the form $(s,t)$ where $s\in S$, $t\in[0,1]$.
The geometric conditions on $M$ will be phrased in terms of “buffer width,” which makes sense in a fairly general setting.
Definition 1.2.
Let $X$ be a geodesic metric space, and let $Y\subset X$ be closed. The buffer width of $Y$ in $X$, written $\operatorname{BW}_{X}(Y)$, is half the length of the shortest non-degenerate local geodesic intersecting $Y$ only in its endpoints.
Example 1.3.
For $N$ a non-positively curved manifold without boundary, and $x\in N$, the buffer width $\operatorname{BW}_{N}(x)$ is equal to the injectivity radius at $x$, usually written $\operatorname{inj}(x)$. If $Y\subset N$ is a totally geodesic submanifold, then $\operatorname{BW}_{X}(Y)$ is the normal injectivity radius of $Y$.
Remark 1.4.
If $\operatorname{BW}_{X}(Y)$ is positive, then $Y$ is locally convex in $X$. See Definition 2.3 for the definition of local convexity.
Here is the first main result. (Recall $\operatorname{inj}(X)=\inf_{x\in X}\operatorname{inj}(x)$.)
Theorem A.
Let $M$ be a hyperbolic manifold with totally geodesic boundary and suppose there are constants $b$ and $c$ so that
(A1) $0<b<\operatorname{BW}_{M}(\partial M),$
and
(A2) $\operatorname{inj}(\partial M)\geq c>\pi/\sinh(b).$
Then there is a negatively curved metric $\hat{d}$ on $\widehat{M}$ and a locally isometric embedding of $M\smallsetminus N_{b}(\partial M)$ into $(\widehat{M},\hat{d})$ with image equal to $M\subset\widehat{M}$.
Remark 1.5.
There exist $b$ and $c$ as in the statement (and thus a negatively curved metric on $\widehat{M}$) whenever
($\clubsuit$) $\operatorname{inj}(\partial M)\sinh(\operatorname{BW}_{M}(\partial M))>\pi.$
If one is only interested in the existence of a negatively curved metric there is therefore a slightly cleaner statement. The extra notation makes it slightly easier to state and prove Theorem B below.
Note that the condition ($\clubsuit$1.5) in dimension $2$ implies an area condition. If $M$ is a hyperbolic surface of genus $g$ with $k$ geodesic boundary circles, these circles will have disjoint collar neighborhoods of area at least $2\operatorname{inj}(\partial M)\sinh(\operatorname{BW}_{M}(\partial M))$. In particular, condition ($\clubsuit$1.5) implies
(1) $\operatorname{Area}(M)>2k\pi.$
Using Gauss-Bonnet, we have $\operatorname{Area}(M)=2\pi(2g+k-2)$. Combining with (1), we see that $g>1$. This is obviously a necessary condition, since $\widehat{M}$ is a closed surface of genus $g$.
Corollary 1.6.
Let $M_{0}$ be a compact hyperbolic $n$–manifold with totally geodesic boundary. There is a finite-sheeted cover $M\to M_{0}$ so that $\widehat{M}$ admits a negatively curved metric.
Proof.
Let $b=\frac{1}{2}\operatorname{BW}_{M_{0}}(\partial M_{0})$, and note that for any finite-sheeted cover $M\to M_{0}$, we have $\operatorname{BW}_{M}(\partial M)=\operatorname{BW}_{M_{0}}(\partial M_{0})$. There are finitely many conjugacy classes $c_{1},\ldots,c_{k}$ in $\pi_{1}M_{0}$ representing nontrivial loops of length at most $2\pi/\sinh(b)$ in $\partial M_{0}$. Since $\pi_{1}M_{0}$ is linear, it is residually finite [Mal40]. In particular there is a finite index normal subgroup of $\pi_{1}M_{0}$ missing the conjugacy classes $c_{1},\ldots,c_{k}$. The corresponding finite-sheeted cover $M\to M_{0}$ satisfies $\operatorname{inj}(\partial M)>\pi/\sinh(b)$. ∎
Note in the above that any further finite-sheeted cover $M^{\prime}\to M$ also has the property that $\widehat{M}^{\prime}$ can be given a negatively curved metric.
Our second main result gives conditions under which a subset $S\subset M$ gives rise to a locally convex (and hence $\pi_{1}$–injective) subset $\widehat{S}\subset\widehat{M}$. We will express these conditions in terms of projections to $\partial M$ as follows.
Definition 1.7 (Projections of levels of $S$ to $\partial M$).
For any $t\in(0,\operatorname{BW}_{M}(\partial M))$, orthogonal projection gives a diffeomorphism $\pi_{t}\colon\thinspace\partial N_{t}(\partial M)\to\partial M$. For such $t$ and any $S\subset M$, we define
$P_{t}(S)=\pi_{t}(S\cap\partial N_{t}(\partial M)).$
For $t=0$, we define $P_{0}(S)=S\cap\partial M$.
Theorem B.
Let $M$ be hyperbolic with totally geodesic boundary, and suppose $S\subset M$ is a closed locally convex set. For $t\in[0,\operatorname{BW}_{M}(\partial M))$, let $P_{t}=P_{t}(S)$, as in Definition 1.7. Suppose that $b,c>0$ satisfy the conditions (A1) and (A2) of Theorem A and additionally that there is a $b^{\prime}\in(b,\operatorname{BW}_{M}(\partial M))$ so that the following hold.
  1. (B1)
    There is an isotopy $\Phi\colon\thinspace[0,b^{\prime}]\times\partial M\to\partial M$ so that $\Phi_{0}$ is the identity and $\Phi_{t}(P_{0})=P_{t}$ for all $t\in[0,b^{\prime}]$.
  2. (B2)
    If $0\leq t<s<b^{\prime}$, then $P_{t}\supseteq P_{s}$, and
  3. (B3)
    $\operatorname{BW}_{\partial M}(P_{t})>\frac{c}{2}$ for all $t\in[0,b^{\prime})$.
Then there is a negatively curved metric on $\widehat{M}$ with respect to which $\widehat{S}$ is locally convex.
Remark 1.8.
The metric in Theorem B can be obtained from the metric in Theorem A via an isotopy.
Remark 1.9.
The hypotheses of Theorem B may be clarified by considering the situation of a properly embedded totally geodesic submanifold $S$ meeting the boundary of $M$. For appropriately chosen $b,b^{\prime}$, $S$ will always satisfy Condition 1. However, $S$ only satisfies hypothesis 2 when it meets the boundary orthogonally, in which case $P_{t}=P_{0}$ for small $t$. (If $S$ does not meet the boundary orthogonally, there may nonetheless be a “thickening” which satisfies hypothesis 2.) Condition 3 can often be ensured in a finite-sheeted cover, as the next Corollary shows.
Corollary 1.10.
Let $M_{0}$ and $b,c>0$ satisfy the conditions of Theorem A, and let $S_{0}\subset M_{0}$ and $b^{\prime}>0$ satisfy all the conditions of Theorem B except possibly Condition 3. Suppose further that $S_{0}$ is connected and $\pi_{1}S_{0}$ is separable in $\pi_{1}M_{0}$. Then there is a finite-sheeted cover $M\to M_{0}$ so that $S$ lifts to $S\subset M$ and a negatively curved metric $\hat{d}$ on $\widehat{M}$ so that $\widehat{S}$ is locally convex in $\widehat{M}$.
Proof.
Scott’s characterization of subgroup separability [Sco78] means that any compact subset $K$ of the cover $\tilde{M}_{0}$ of $M_{0}$ corresponding to $\pi_{1}S_{0}$ embeds in a finite-sheeted cover $M$ of $M_{0}$. The locally convex set $S_{0}$ lifts to a set $\tilde{S}_{0}$ in $\tilde{M}_{0}$ with infinite buffer width. Taking $K$ to be a closed $c$–neighborhood of $\tilde{S}_{0}$, the corresponding finite cover $M$ contains a lifted copy $S$ of $S_{0}$, which moreover satisfies condition 3. ∎
We remark that the condition of separability is always true if $n\leq 3$ by results of Scott [Sco78] and Wise [Wis21].
We state a group-theoretic corollary.
Corollary 1.11.
Suppose that $M$ is a compact hyperbolic $n$–manifold with totally geodesic boundary satisfying the conditions of Theorem A. Let $\Gamma=\pi_{1}M$, and let $K\lhd\Gamma$ be the subgroup normally generated by loops in $\partial M$. Then $\Gamma/K$ is torsion-free, hyperbolic, and has cohomological dimension $n$.
Suppose in addition that $H<\Gamma$ is the fundamental group of a locally convex subset $S\subset M$ satisfying the conditions of Theorem B. Then the image of $H$ in $\Gamma/K$ is quasi-convex.
Proof.
The first part of the Corollary is immediate from Theorem A. If each component of $\partial M$ intersects $S$ either in a connected set or in the empty set, the second part is immediate from Theorem B, since the image $\bar{H}$ of $H$ in $\Gamma/K$ can be identified with the fundamental group of $\widehat{S}$. Since $\widehat{M}$ is negatively curved and $\widehat{S}$ is a locally convex subset, the fundamental group of $\widehat{S}$ injects into that of $\widehat{M}$ (see [BH99, II.4.14]). If there are components of $\partial M$ whose intersection with $S$ is disconnected, then $\pi_{1}\widehat{S}$ is still a quasi-convex subgroup, but is not equal to $\bar{H}$; rather, it contains $\bar{H}$ as a free factor. But a free factor of a group is always quasi-isometrically embedded, so the quasi-convexity of $\bar{H}$ follows from the quasi-convexity of $\pi_{1}\widehat{S}$. ∎
The first part of Corollary 1.11 is related to the Relatively Hyperbolic Dehn Filling Theorem [GM08, Osi07] – the statement about cohomological dimension is related to work of Sun and Petrosyan [PS24]. The second part is connected to [GM21, Proposition 2.5]. All those statements though are for “sufficiently large” or “sufficiently $H$–wide” fillings. In principle the bounds for those results can be made explicit, but will not be as tight as the ones which we have provided.
Acknowledgments
We would like to thank Daniel Groves for helpful conversations related to the correct statement of Theorem B. CK was supported by the NSF under grant DGE-1650441. JM was partially supported by the Simons Foundation, grant number 942496.
2. Preliminaries
2.1. Warped products
We recall here the definition of a warped product of length spaces, which generalizes a warped product of Riemannian manifolds. For more information see [BBI01, §3.6.4].
Definition 2.1.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be complete length spaces and let $f\colon\thinspace X\to[0,\infty)$ be a continuous function. For a lipschitz path $\gamma\colon\thinspace[a,b]\to X\times Y$, write $\gamma_{X}$ and $\gamma_{Y}$ for the projections to $X$ and $Y$ respectively. These are also lipschitz paths so their speeds $|\gamma_{X}^{\prime}|$, $|\gamma_{Y}^{\prime}|$ are defined almost everywhere [BBI01, Theorem 2.7.6]. Define a length
[…]
and use this to define a length pseudo-metric $d_{f}$. The warped product $X\times_{f}Y$ is the canonical metric space quotient of $(X\times Y,d_{f})$. (This is homeomorphic to $X\times Y$ if $f$ has no zeroes; otherwise some copies of $Y$ are smashed to points.)
If $(X,g_{X})$ and $(Y,g_{Y})$ are Riemannian manifolds, and $f$ has no zeroes, then the warped product metric on $X\times_{f}Y$ is induced by the Riemannian metric
$g=g_{X}+f^{2}g_{Y}.$
For example, the warping function $f\equiv 1$ yields the standard product metric. More interestingly, hyperbolic space $\mathbb{H}^{n+1}$ is isometric to a warped product in many ways, for example ×$\mathbb{R}~{}\times_{e^{t}}~{}\mathbb{E}^{n}$ or $[0,\infty)\times_{\sinh(t)}\mathbb{S}^{n}$. Most relevant for our purposes is the isometry
$\mathbb{H}^{n+1}\cong\mathbb{R}\times_{\cosh(t)}\mathbb{H}^{n}.$
We record the following consequence for later use.
Remark 2.2.
For a compact hyperbolic manifold $M$ with totally geodesic boundary, and $r\leq\operatorname{BW}_{M}(\partial M)$, there is a Riemannian isometry
$N_{r}(\partial M)\cong[0,r)\times_{\cosh(t)}\partial M.$
The right-hand side can be identified with a subset of the normal bundle of $\partial M$ in $M$, endowed with the pullback of the hyperbolic metric on $M$ under the normal exponential map.
2.2. Local convexity
Definition 2.3.
Let $Z$ be a geodesic space and $S\subseteq Z$ a subset. The set $S$ is convex if every geodesic in $Z$ with endpoints in $S$ lies in $S$.
The set $S$ is locally convex if for every point $s\in S$ there is a neighborhood $U$ of $s$ in $Z$ so that $S\cap U$ is convex in $Z$.
Definition 2.4.
The local convexity radius at a point $s$ of a subset $S$ of a geodesic space $Z$ is
$\mathrm{lcr}_{S}(s)=\sup\{r\in\mathbb{R}:B_{r}(s)\cap S\mbox{ is convex in Z}\}.$
The local convexity radius of $Q\subseteq S$ in $Z$ is
$\mathrm{lcr}_{S}(Q)=\inf_{s\in Q}\mathrm{lcr}_{S}(s).$
Lemma 2.5.
Let $Z$ be a complete non-positively curved metric space, and $S$ a compact, locally convex subset of $Z$. The local convexity radius of $S$ in $Z$ is positive.
Proof.
We first show the function $\mathrm{lcr}_{S}:S\to\mathbb{R}$ is lower semi-continuous. Fix $s\in S$ and $\varepsilon>0.$ Without loss of generality, we may assume that
$\varepsilon<\min\{\operatorname{BW}_{Z}(s),\mathrm{lcr}_{S}(s)\}.$
We must show that there exists $\delta>0$ such that $\mathrm{lcr}_{S}(s)-\varepsilon<\mathrm{lcr}_{S}(c)$ for all $c$ in $B_{\delta}(s)\cap S$. Set $\delta=\frac{\varepsilon}{2}.$ Let $c$ be a point in $B_{\delta}(s)\cap S$. Any ball $B_{r}(c)$ which is contained in $B_{\mathrm{lcr}_{S}(s)}(s)$ will have convex intersection with $S$; this can be seen by lifting to the universal cover of $Z$ (which we may do because $\varepsilon<\operatorname{BW}_{Z}(s)$). In particular, since
$\delta=\frac{\varepsilon}{2}<\mathrm{lcr}_{S}(s)-d_{Z}(s,c),$
the intersection $B_{\delta}(c)\cap B_{\mathrm{lcr}_{S}(s)}(s)$ is convex in $Z$. The intersection of two convex sets is convex, so for such $r$, $B_{\mathrm{lcr}_{S}(s)}(s)\cap S\cap B_{\delta}(c)=S\cap B_{\delta}(c)$ is convex.
Now suppose by way of contradiction that there exists a sequence of points $(s_{i})\subset S$ such that $\mathrm{lcr}_{S}(\{s_{i}\})=0$. Since $S$ is compact, we may suppose that $(s_{i})$ converges to some $s_{\infty}\in S$. Since $\mathrm{lcr}_{S}:S\to\mathbb{R}$ is lower semi-continuous,
$0=\lim_{i\to\infty}\mathrm{lcr}_{S}(s_{i})\geq\mathrm{lcr}_{S}(s_{\infty}),$
but $\mathrm{lcr}_{S}(s_{\infty})>0$ since $s_{\infty}$ is in the locally convex set $S$. ∎
3. Tailoring warped products to bound curvature from above
Our first goal is to construct a warping function $f$ on an interval $I\subset\mathbb{R}$ such that $f$ vanishes at the left-hand endpoint of $I$, $f$ makes the metric on the cone agree with the hyperbolic metric on $M$ away from a neighborhood of the cone point, and the cone $I\times_{f}N$ is $\operatorname{CAT}(K)$ for some $K<0$, where $N$ is a boundary component of $M$. We will use a criterion of Alexander and Bishop [AB04] to bound the curvature of our warped product from above. They work in a more general setting $B\times_{f}F$ where $B$ need not be an interval, so we simplify their definitions and result here.
Definition 3.1.
[AB04, §2.2] Let $\mathcal{F}K$ denote the family of solutions to the differential equation $g^{\prime\prime}+Kg=0$. An $\mathcal{F}K$-convex function $f:I\to\mathbb{R}$ for $I\subseteq\mathbb{R}$ an interval is one whose restriction to every sub-interval satisfies the differential inequality $f^{\prime\prime}\geq-Kf$ in the barrier sense, meaning if $g\in\mathcal{F}K$ coincides with $f$ at the endpoints of a sub-interval, then $f\leq g$.
Theorem 3.2.
[AB04, Theorem 1.1] Let $I=[t_{0},t_{1})\subset\mathbb{R}$ and let $F$ be a complete $\operatorname{CAT}(K_{F})$ space for some $K_{F}\in\mathbb{R}$. Fix some $K\leq 0$ and suppose $f:I\to[0,\infty)$ is $\mathcal{F}K$-convex, with $f^{-1}(0)=\{t_{0}\}$. Suppose further that $f^{\prime}({t_{0}}^{+})^{2}\geq K_{F}$.
Then the warped product $I\times_{f}F$ is $CAT(K)$.
The following lemma gives us sufficient conditions for such an $f$ to be $\mathcal{F}K$-convex.
Lemma 3.3.
Suppose $K<0$ and $I\subseteq\mathbb{R}$ is an interval. Let $f:I\to\mathbb{R}$ be a function with first derivative defined everywhere and second derivative defined almost everywhere. If $f$ satisfies the differential inequality
$f^{\prime\prime}+Kf\geq 0$
wherever $f^{\prime\prime}$ is defined, then $f$ is $\mathcal{F}K$-convex.
Here is a lemma that will help us construct a warping function that we can use when applying Theorem 3.2 of Alexander-Bishop.
Lemma 3.4.
For all $b>0$ and $0<\delta<\sinh(b)$, there exists $t_{0}<b$ and a function $f\colon\thinspace[t_{0},\infty)\to\mathbb{R}$ such that
  1. (1)
    $f(t_{0})=0$ and $f(t)=\cosh(t)$ for $t\geq b$,
  2. (2)
    $f^{\prime}({t_{0}}^{+})=\delta$, and $f^{\prime}$ is continuous and positive everywhere,
  3. (3)
    $f$ is $\mathcal{F}K$-convex for some $K<0$.
Definition 3.5.
Call a function satisfying the conclusions of Lemma 3.4 a cone warping function with parameters $(b,\delta,K)$.
Proof.
Fix $b>0$ and $\delta\in(0,\sinh(b))$. Let $l_{0}$ be the tangent line to $\cosh(t)$ at $t=b$, i.e.
$l_{0}(t)=\sinh(b)(t-b)+\cosh(b).$
Choose $t_{0}<b$ so that the line of slope $\delta$ through the point $(0,t_{0})$ intersects $l_{0}$ somewhere on the interval $(t_{0},b)$. Let $l_{1}$ be this second line. By [FM10, Lemma 4.13], there is a smooth function $g$ on $[t_{0},b]$ whose graph is tangent to $l_{0}$ at $(b,\cosh(b))$ and to $l_{1}$ at $(t_{0},b)$, and whose second derivative is bounded below everywhere. In particular we have the following, for some $\mu>0$.
  1. (1)
    $g(t_{0})=0$ and $g(b)=\cosh(b)$;
  2. (2)
    $g^{\prime}(t_{0})=\delta$ and $g^{\prime}(b)=\sinh(b)$;
  3. (3)
    $g^{\prime\prime}(t)\geq\mu$ for all $t\in[t_{0},b]$
Define our desired function $f$ as follows:
$f(t):=\begin{cases}g(t)&t\in[t_{0},b]\\ cosh(t)&t\geq b.\end{cases}$
Figure 1. An example $f(t)$ in pink and purple solid lines with tangent lines in dashed green and orange.
See Figure 1 for an example graph of $f$. Note that items (1) and (2) have been satisfied by construction. It remains to show that $f$ is $\mathcal{F}K$-convex for some $K<0$. By Lemma 3.3, it is enough to find $K<0$ so that $f^{\prime\prime}(t)+Kf(t)\geq 0$ for all $t\in(0,b)\cup(b,\infty)$. For $t\in(b,\infty)$, $f^{\prime\prime}(t)=f(t)$, so any $K\geq-1$ works for this interval. For $t\in(t_{0},b)$, we have $f(t)<\cosh(b)$ and $f^{\prime\prime}\geq\mu$, so
$f^{\prime\prime}(t)>\frac{\mu}{\cosh(b)}f(t).$
Taking $K=\max\{-1,-\frac{\mu}{\cosh(b)}\}$, we have $f^{\prime\prime}(t)+Kf(t)\geq 0$ for all $t\in(t_{0},b)\cup(b,\infty)$. Lemma 3.3 gives us conclusion (3). ∎
To apply Alexander–Bishop’s Theorem 3.2 with a component of $\partial M$ as a fiber, we need to find a $\kappa>0$ so that $\partial M$ is $\operatorname{CAT}(\kappa)$. The following finds such a $\kappa$ in terms of the injectivity radius.
Lemma 3.6.
Let $N$ be a complete connected non-positively curved manifold, and let
$\kappa=\left(\frac{\pi}{\operatorname{inj}(N)}\right)^{2}.$
Then $N$ is $\operatorname{CAT}(\kappa)$.
Proof.
Recall that the $\operatorname{CAT}(\kappa)$ inequality concerns only geodesic triangles with perimeter less than $2D_{\kappa}$, where $D_{\kappa}=\frac{\pi}{\sqrt{\kappa}}$ is the diameter of the corresponding model space. Let $\Delta$ be a geodesic triangle in $N$ with perimeter less than $2D_{\kappa}$ and vertices labeled $a,b,c$. Because of the triangle inequality, the length of each side of $\Delta$ is less than $\frac{\pi}{\sqrt{\kappa}}$. Let $x$ and $y$ be two points on $\Delta$ not on the same side. Relabeling vertices if necessary, we may assume $x\in[a,b]$ and $y\in[a,c]$. Since the sides $[a,c]$ and $[a,b]$ are of length less than the injectivity radius of $N$, the points $x$ and $y$ are in an open ball $B$ around $a$ of radius less than $\operatorname{inj}(N)$. We have that
(2) $d_{N}(x,y)\leq d_{B}(x,y),$
and we can lift $B$ to a ball in the universal cover $\widetilde{N}$. Since balls in $\widetilde{N}$ are convex, it follows that
$d_{B}(x,y)=d_{\widetilde{N}}(x,y).$
Since the sectional curvature of $N$ is bounded above by 00, so too is the sectional curvature of $\widetilde{N}$. Thus $\widetilde{N}$ is $\operatorname{CAT}(0)$ (cf. Theorem II.1A.6 in [BH99]), which implies it is $\operatorname{CAT}(\kappa)$. The ball $B$ embeds isometrically so $d_{B}(x,y)$ is bounded above by the corresponding distance in the $\operatorname{CAT}(\kappa)$ comparison triangle. By (2) this also gives an upper bound for $d_{N}(x,y)$, verifying the $\operatorname{CAT}(\kappa)$ inequality. ∎
4. Proof of Theorem A
We assume for this section and the next that the hyperbolic manifold $M$ with totally geodesic boundary satisfies the hypotheses of Theorem A. Namely, there are constants $b,c$ satisfying
(A1) $0<b<\operatorname{BW}_{M}(\partial M),$
and
(A2) $\operatorname{inj}(\partial M)\geq c>\pi/\sinh(b).$
For the rest of the section we fix these constants and also fix
$\delta=\frac{\pi}{c},\quad t_{0}<b,\quad K<0$
and a cone warping function $f\colon\thinspace[t_{0},\infty)\to\mathbb{R}$ with parameters $(b,\delta,K)$. Since $\delta<\sinh(b)$ such a function exists by Lemma 3.4.
Lemma 4.1.
Let $N$ be a boundary component of $M$. There exists a function $f$ such that the warped cone [^1] We use the notation $C_{f}(N)$ to distinguish from $C(N)=N\times[0,1]/(N\times\{0\})$, the topological cone.
$C_{f}(N)=[t_{0},\operatorname{BW}_{M}(\partial M))\times_{f}N$
is $\operatorname{CAT}(K)$ for some $K<0$. Moreover the subset $[b,\operatorname{BW}_{M}(\partial M))\times_{f}N$ is (Riemannian) isometric to $N_{\operatorname{BW}_{M}(\partial M)}(N)\smallsetminus N_{b}(N)$.
Proof.
Assumption (A2) together with Lemma 3.6 implies that $N$ is $\operatorname{CAT}(K_{F})$ where
$K_{F}=\left(\frac{\pi}{c}\right)^{2}=\delta^{2}.$
Item (2) of Lemma 3.4 implies our cone warping function $f$ satisfies $f^{\prime}({t_{0}}^{+})=\delta$. Item (3) implies that $f$ is $\mathcal{F}K$–convex, so the Alexander–Bishop theorem 3.2 implies that the cone $C_{f}(N)$ is $\operatorname{CAT}(K)$.
By item (1), $f(t)=\cosh(t)$ for all $t\in[b,\operatorname{BW}_{M}(\partial M))$. As noted in Remark 2.2, the $\operatorname{BW}_{M}(\partial M)$–neighborhood of $\partial M$ is a warped product with warping function $\cosh(t)$. In particular, the subset
$[b,\operatorname{BW}_{M}(\partial M))\times_{f}N\subset C_{f}(N)$
is Riemannian isometric to
$N_{\operatorname{BW}_{M}(\partial M)}(N)\smallsetminus N_{b}(N).$
See Figure 2
Figure 2. On the left is a collar neighborhood of width less than $\operatorname{BW}_{M}(\partial M)$ of a component $N$ of $\partial M$. On the right is the corresponding cone in $X$. (Note that although the cone angle is small in the cartoon it will be at least $2\pi$.)
for a schematic. ∎
We will use these warped cones to describe a geodesic space $X$ homeomorphic to $\widehat{M}$. Let $\mathcal{N}$ denote the set of boundary components of $M$. For each $N\in\mathcal{N}$ there is a normal exponential map
$E_{N}\colon\thinspace N\times[0,\operatorname{BW}_{M}(\partial M))\to M,$
so that, for fixed $x$, the map $t\mapsto E_{N}(x,t)$ gives a unit speed geodesic in $M$, orthogonal to $N$ and starting at $x$.
Definition 4.2 (Our metric model for $\widehat{M}$).
Let $\overline{M}=M\smallsetminus N_{b}(\partial M)$. We define
($\heartsuit$) $X=\overline{M}\sqcup\left(\bigsqcup\{C_{f}(N)\mid N\in\mathcal{N}\}\right)/\sim$
where $\sim$ is the equivalence relation generated by the gluings
$E_{N}(x,t)\sim(t,x)\in C_{f}(N),\quad N\in\mathcal{N},\ x\in N,\ t\in[b,\operatorname{BW}_{M}(\partial M)).$
The metrics on the pieces $C_{f}(N)$ and $\overline{M}$ together induce a path metric on $X$.
Lemma 4.3.
The path metric space $X$ defined by (italic-$\heartsuit$4.2) is negatively curved.
Proof.
Note that because $f(t)=\cosh(t)$ on $[b,\operatorname{BW}_{M}(\partial M))$, the gluings are by (Riemannian) isometries. Moreover each piece gives an open set in $X$ which is either locally $\operatorname{CAT}(-1)$ (in the case of $\overline{M}$) or $\operatorname{CAT}(K)$ for some $K<0$ (in the case of the $C_{f}(N)$ pieces). In particular the space $X$ is negatively curved. ∎
Recall that the space $\widehat{M}$ was defined in Definition 1.1 as a quotient of $M\sqcup(\partial M\times[0,1])$ where $\partial M\times\{1\}$ is identified with $\partial M\subset M$, and each component of $\partial M\times\{0\}$ is collapsed to a point. The space $X$ is defined in Definition 4.2.
Lemma 4.4.
There is a homeomorphism
$\phi\colon\thinspace X\to\widehat{M}$
with the following properties:
  1. (1)
    The image of $\overline{M}$ in $\widehat{M}$ is $M$.
  2. (2)
    If $d_{M}(x,\partial M)\geq\operatorname{BW}_{M}(\partial M)$, then $\phi(x)=x$.
  3. (3)
    For each boundary component $N$ of $M$, the image of $C_{f}(N)$ in $\widehat{M}$ is the cone $N\times[0,1]/\sim$.
  4. (4)
    There is a diffeomorphism $\xi\colon\thinspace[b,\operatorname{BW}_{M}(\partial M))\to[0,\operatorname{BW}_{M}(\partial M))$ so that, for each component $N$ of $\partial M$, each $x\in N$, and each $t\in[b,\operatorname{BW}_{M}(\partial M)$, we have
    $\phi(E_{N}(x,t))=E_{N}(x,\xi(t)).$
Proof.
We exhibit a homeomorphism $\phi$ from $X$ to $\widehat{M}$ by defining maps $\phi_{M}\colon\thinspace\overline{M}\to\widehat{M}$ and $\phi_{C}\colon\thinspace C_{f}(N)\to\widehat{M}$ on the pieces from the definition ($\heartsuit$4.2) of $X$, and showing they agree on the overlaps.
To begin, we choose a diffeomorphism $\xi\colon\thinspace[b,\operatorname{BW}_{M}(\partial M))\to[0,\operatorname{BW}_{M}(\partial M))$. Let $E_{\partial M}(x,t)=E_{N}(x,t)$ if $x$ is in the component $N$ of $\partial M$ and $t\in[0,\operatorname{BW}_{M}(\partial M))$. For the orthogonal projection map $\pi_{t}$ defined in Definition 1.7, and for any such $x,t$, we have
$\pi_{t}(E_{\partial M}(x,t))=x.$
Now we define $h\colon\thinspace\overline{M}\to M$ by using $\xi$ to stretch a regular neighborhood of $\partial\overline{M}$ to fit onto a regular neighborhood of $\partial M$:
$h(x)=\begin{cases}x&x\in M\smallsetminus N_{\operatorname{BW}_{M}(\partial M)}(\partial M)\\ E_{\partial M}(\pi_{t}(x),\xi(t))&x\in E_{\partial M}(\partial M,t),\ t\in[0,\operatorname{BW}_{M}(\partial M))\end{cases}.$
This is a homeomorphism because of assumption (A1). If $\iota_{M}\colon\thinspace M\to\widehat{M}$ is the inclusion, we set
$\phi_{M}=\iota_{M}\circ h.$
Note that for any boundary component $N$ of $M$, the map $\phi_{M}$ already determines $\phi_{C}(t,x)$ for $t\in[b,\operatorname{BW}_{M}(\partial M))$; we next need to extend $\phi_{C}$ to the rest of the cone. To do this, let $\alpha\colon\thinspace[t_{0},b]\to[0,1]$ be an orientation preserving homeomorphism. For $t\in[t_{0},b]$ and $x\in N$, define
$\phi_{C}(t,x)=(x,\alpha(t)),$
[…]
Proof of Theorem A.
Lemma 4.3 shows that the space $X$ is negatively curved. Lemma 4.4 shows that $X$ is homeomorphic to $\widehat{M}$. To obtain the metric $\hat{d}$ on $\widehat{M}$, we push forward the path metric on $X$.
The desired local isometric embedding of $\overline{M}$ into $\widehat{M}$ comes from item (1) in Lemma 4.4. ∎
5. Proof of Theorem B
In this section we retain the assumptions made at the beginning of the last section, in particular fixing the constants $b,c,\delta,t_{0},K$, and the cone warping function $f$ with parameters $(b,\delta,K)$. We moreover fix is a closed locally convex subset $S$ in $M$ and a $b^{\prime}\in(b,\operatorname{BW}_{M}(\partial M))$ so the additional assumptions of Theorem B all hold.
Our model $X$ for $\widehat{M}$ is described above in Definition 4.2. It consists of a constant curvature part, metrized as $\overline{M}=M\smallsetminus N_{b}(\partial(M))$, together with the warped cones
$C_{f}(N)=[t_{0},\operatorname{BW}_{M}(\partial M))\times_{f}N$
defined for each boundary component $N$ of $M$. The metric on the union comes from gluing together the path metrics on the pieces.
We now define a subset of $X$ which will be our model for $\widehat{S}$. Namely, let
$Y=(S\smallsetminus N_{b}(\partial M))\cup\left(\bigsqcup_{N\in\mathcal{N}}\{(t,s)\in C_{f}(N):s\in P_{b},t\in[t_{0},b]\}\right).$
Figure 3. A cartoon of $S\subset M$ (on the left) and $Y\subset X$ (on the right). The regions above level $b$ are isometric. Again, the cone angle on the right is pictured as small, but in the construction will be large.
See Figure 3 for a schematic.
We must show that $Y$ is locally convex in $X$. This will follow from three lemmas.
Lemma 5.1.
$Y$ is locally convex at any point whose distance from the cone point set is greater than $b-t_{0}$.
Lemma 5.2.
$Y$ is locally convex at any non-cone-point in $Y\cap(\bigsqcup C_{f}(N))$.
Lemma 5.3.
$Y$ is locally convex at any cone point.
The lemmas will be proved after the proof of Theorem B.
Proof of Theorem B.
Lemmas 5.1, 5.2, and 5.3 account for all points in $Y$, so $Y$ is locally convex in $X$.
Lemma 4.4 gives a homeomorphism
$\phi\colon\thinspace X\to\widehat{M}=M\cup(\partial M\times[0,1])/\sim,$
which sends each point $x$ of $\overline{M}$ in $X$ either to $x$ (if it is outside the $\operatorname{BW}_{M}(\partial M)$–neighborhood of the boundary), or to a point on a shortest path from $x$ to $\partial M$. Points inside the $\operatorname{BW}_{M}(\partial M)$–neighborhood of $\partial M$ which are equidistant from the boundary are sent to points which are equidistant from the boundary.
Our assumption 1 together with Lemma 4.4 will ensure that the homeomorphism $\phi$ sends $Y$ to a set which is isotopic to $\widehat{S}$. Specifically, we use the isotopy $\Phi\colon\thinspace[0,b^{\prime}]\times\partial M\to\partial M$ to define an isotopy $\Psi\colon\thinspace[0,1]\times\widehat{M}\to\widehat{M}$ as follows:
  • On $M\smallsetminus N_{b^{\prime}}(\partial M)$: Define $\Psi_{s}(x)=x$ for all $s\in[0,1]$.
  • On $N_{b^{\prime}}(\partial M)$: Recall that we may identify $N_{b^{\prime}}(\partial M)$ with $[0,b^{\prime})\times_{\cosh t}\partial M$. Using the coordinates from this identification, define
    $\Psi_{s}(t,x)=(t,\Phi_{st}\circ(\Phi_{s\xi^{-1}(t)})^{-1}(x)).$
  • On a cone $N\times[0,1]/(N\times\{0\})$: Define
    $\Psi_{s}(x,t)=(\Phi_{sb}^{-1}(x),t).$
The reader may check that this gives a well-defined isotopy taking $\phi(Y)$ to $\widehat{S}$. ∎
It remains to prove the lemmas. Here is some notation that will be helpful. For a point $x$ in some $C_{f}(N)\subset X$, we can write $x=(t(x),p(x))$ where $t(x)$ is in $[t_{0},\operatorname{BW}_{M}(\partial M))$ and $p(x)$ is in $N$. (For $t(x)\in[b,\operatorname{BW}_{M}(\partial M))$ such a point also corresponds to a point in $\overline{M}$, but this ambiguity will cause no problems.)
Proof of Lemma 5.1.
Suppose that $y\in Y$ is farther than $b-t_{0}$ from any cone point. Then $y$ is in the interior of $\overline{M}$. In particular there is an $\varepsilon>0$ so that the pair $(B_{\varepsilon}(y),B_{\varepsilon}(y)\cap Y)$ in $X$ is isometric to the pair $(B_{\varepsilon}(y),B_{\varepsilon}(y)\cap S)$ in $M$. The local convexity radius of $Y$ at $y$ in $X$ is therefore at least the minimum of $\varepsilon$ and $\mathrm{lcr}_{S}(S)$, which is positive by Lemma 2.5. ∎
Proof of Lemma 5.2.
Fix $y=(t(y),p(y))$ in $Y\cap C_{f}(N)$ for some boundary component $N$. If $t(y)>b$, then $Y$ is locally convex at $y$ by Lemma 5.1, so we suppose $t(y)\leq b$. Let
$L=\frac{2}{3}(t(y)-t_{0})+t_{0}.$
(This is $\frac{2}{3}$ the distance from the cone point to $y$.) We set
(3) $\varepsilon<\min\left\{\frac{b^{\prime}-t(y)}{3},\frac{L}{2},\frac{cf(L)}{2},\frac{\mathrm{lcr}_{S}(S)}{2}\right\}.$
Note that all four of the quantities on the right hand side of (3) are positive; $\mathrm{lcr}_{S}(S)/2$ by Lemma 2.5, and the others because $t(y)\in(t_{0},b^{\prime})$.
We first show that $B_{\varepsilon}(y)$ is convex in $M$. Let $z,w\in B_{\varepsilon}(y)$. By the triangle inequality, any shortest path $[z,w]$ stays inside $B_{3\varepsilon}(y)$. Since $\varepsilon<\frac{1}{3}(b^{\prime}-t(y))$, $B_{3\varepsilon}(y)$ is contained in $C_{f}(N)$. In particular any shortest path in $X$ joining $z$ to $w$ is also a shortest path in the warped product metric on $C_{f}(N)$ (as opposed to a path leaving $C_{f}(N)$). By Lemma 4.1, $C_{f}(N)$ is $\operatorname{CAT}(0)$. In particular, balls in $C_{f}(N)$ are convex, so the segment $[z,w]$ lies inside $B_{\varepsilon}(y)$ as desired. This argument also shows
$d_{X}(z,w)=d_{C_{f}(N)}(z,w).$
We now assume $z,w$ are in $Y\cap B_{\varepsilon}(y)$. Let $u$ be a point on $[z,w]$. We must show $u\in Y$. Because the warping function $f$ is increasing, we always have $t(u)\leq\max\{t(z),t(w)\}$. To show $u\in Y$, we must show either that $p(u)\in P_{t(u)}$ (if $t(u)>b$), or that $p(u)\in P_{b}$ (if $t(u)\leq b$). To simplify, we introduce the following notation
$P^{*}_{t}=\begin{cases}P_{t}&t>b\\ P_{b}&t\leq b.\end{cases}$
In this language we must show $p(u)\in P^{*}_{t(u)}$.
We divide into cases depending on the relationship between the heights $t(u)$, $t(z)$ and $t(w)$.
Case 1.
The height $t(u)$ is at most $\min\{t(z),t(w)\}$.
We have $P_{t(u)}\supseteq P_{t(z)}\cup P_{t(w)}$ by hypothesis 2, and so $P^{*}_{t(u)}\supseteq P^{*}_{t(z)}\cup P^{*}_{t(w)}$. Notably this means that the endpoints of the geodesic $[p(z),p(w)]$ are in $P^{*}_{t(u)}$. Since $z$ and $w$ are in $B_{\varepsilon}(y)$,
($\diamondsuit$) $d_{X}(z,w)<2\varepsilon.$
Because $\varepsilon<\frac{1}{2}L=\frac{1}{3}(t(y)-t_{0})$, the ball $B_{\varepsilon}(y)$ is disjoint from $[t_{0},L]\times N\subset C_{f}(N)$. Since the warping function $f$ is increasing and $B_{\varepsilon}(y)$ is entirely above $\{L\}\times N$,
($\spadesuit$) $d_{X}(z,w)\geq d_{\{L\}\times N}((L,p(z)),(L,p(w)))=f(L)d_{\partial M}(p(z),p(w)).$
Combining equations ($\diamondsuit$5) and ($\spadesuit$5) yields
$d_{\partial M}(p(z),p(w))<\frac{2\varepsilon}{f(L)}.$
From
$\varepsilon<\frac{cf(L)}{2},$
we deduce
$d_{\partial M}(p(z),p(w))<c.$
The buffer width of $P^{*}_{t(u)}$ in $\partial M$ is assumed in hypothesis 3 to be greater than $\frac{c}{2}$, so $[p(z),p(w)]$ must be contained in $P^{*}_{t(u)}$. In particular $p(u)\in P^{*}_{t(u)}$, as desired.
Case 2.
The height $t(u)$ is between $t(z)$ and $t(w)$.
Without loss of generality suppose $t(w)<t(u)<t(z)$. By the argument in Case 1, $[p(z),p(w)]\subseteq P^{*}_{t(w)}$. For the sake of contradiction, suppose that $u\in[z,w]$ is not in $Y$. Re-choosing $w$ if necessary, we may assume $t(\cdot)$ is monotone on $[z,w]$. Possibly shortening the segment again, we may assume that $[z,w]$ meets $Y$ only in its endpoints. Thus, for $t\in(t(w),t(z))$, we must have ${P}^{*}_{t}\neq P^{*}_{t(w)}$. Hence $t(w)\geq b$, since $P^{*}_{t}=P_{b}$ for all $t\leq b$. This means that $[z,w]$ is actually a shortest path in $M$, contained in some $2\varepsilon$–ball around a point in $S$. Similarly, the points $z,w$ are points of $S$. Since $\varepsilon<\frac{1}{2}\mathrm{lcr}_{S}(S)$, all of $[z,w]$ is contained in $S$, and hence in $Y$. In particular, $u\in Y$, a contradiction. ∎
Proof of Lemma 5.3.
Let $y$ be the cone point of $C_{f}(N)$ for a component $N$ of $\partial M$ so that $N\cap S\neq\emptyset$. Let $\varepsilon=\frac{1}{3}(b-t_{0}).$ For any $z,w\in B_{\varepsilon}(y)$, the triangle inequality shows that a shortest path lies entirely in $C_{f}(N)$, and
$d_{X}(z,w)=d_{C_{f}(N)}(z,w).$
Lemma 4.1 shows $C_{f}(N)$ is $\operatorname{CAT}(0)$, so in fact there is a unique shortest segment $[z,w]$, which lies in $B_{\varepsilon}(y)$.
We suppose now that $z,w$ lie in $Y\cap B_{\varepsilon}(y)$. If $[z,w]$ goes through the cone point it is clearly contained in $Y$, so we assume it avoids the cone point. We can thus write $[z,w]$ as a path
$\sigma(s)=(\tau(s),\rho(s)),\ s\in[0,1],$
where the path $\rho(s)$ is a (not necessarily constant speed) geodesic in $\partial M$. We must show that $\rho(s)\in P_{b}$ for all $s$. By way of contradiction suppose some $\rho(s)\not\in P_{b}$. Replacing $[z,w]$ by a sub-segment if necessary, we may assume that $\rho(s)\not\in P_{b}$ for all $s\in(0,1)$. We therefore have
$d_{\partial M}(\rho(0),\rho(1))\geq 2\operatorname{BW}_{\partial M}(P_{b})>c,$
by Assumption 3. Let $J$ be the interval $[0,d_{\partial M}(\rho(0),\rho(1))]$. The path $\sigma$ has image in a subset of $C_{f}(N)$ isometric to
$W=[t_{0},b)\times_{f}J.$
Since $f$ is convex, the metric in $W$ dominates the metric in the Euclidean cone
$W^{\prime}=[t_{0},b)\times_{\delta(t-t_{0})}J,$
where we recall that
$\delta=f^{\prime}({t_{0}}^{+})=\frac{\pi}{c}.$
In particular,
$d_{M}(z,w)$ $=d_{W}((\tau(0),0),(\tau(1),|J|)$
$\geq d_{W^{\prime}}((\tau(0),0),(\tau(1),|J|)).$
Also note that $W^{\prime}$ is isometric to the standard Euclidean cone
$W^{\prime\prime}=[0,b-t_{0})\times_{t}\delta J.$
Note that $|J|>c$, so the interval $\delta J$ has length greater than $\pi$. In particular the geodesic in $W^{\prime\prime}$ from $(\tau(0)-t_{0},0)$ to $(\tau(1)-t_{0},\delta|J|)$ goes through the cone point. Thus
$d_{M}(z,w)\geq(\tau(0)-t_{0})+(\tau(1)-t_{0}).$
There is a geodesic in $C_{f}(N)$ of length exactly $(\tau(0)-t_{0})+(\tau(1)-t_{0})$, but it goes through the cone point, contradicting our assumption that it doesn’t. This contradiction allows us to conclude that $\rho(s)\in P_{b}$, as desired.